# Simulations of Ellipsoidal Primordial Black Hole Formation

### v2

@@ 1 Introduction
Primordial black holes (PBHs), if they exist, are black holes that could have been formed in the early Universe [1, 2, 3, 4, 5] (see [6, 7, 8, 9, 10, 11] for reviews covering different perspectives). A PBH results from the collapse of a sufficiently large over-density in the early universe, much before the moment of matter-radiation equality, and therefore, their existence can be an evidence of the existence of primordial inhomogeneities. Interestingly, PBHs may constitute a significant fraction of the dark matter [12] (especially in the so-called asteroid mass range) and explain different cosmic conundra [13]. Currently, PBHs have not been detected so far, but future gravitational wave observations may establish their existence (in particular if their mass is lower than a solar mass)[14, 15, 16] and quantify their role in the dark matter.
@@
To accurately estimate the abundance of PBHs in our Universe, it is essential to clarify the initial conditions that lead to their formation. Particularly, determinating the threshold is essential for the abundance estimation as it is exponentially sensitive to the threshold [4]. The threshold estimation typically requires relativistic numerical simulations. Most studies on simulation of PBH formation have focused on spherically symmetric systems (see [17] for a comprehensive review and references therein) for what we have nowadays a relatively solid understanding. In contrast, our understanding of PBH formation in non-spherical scenarios currently remains limited, with only a few studies numerically addressing this issue [18, 19, 20, 21] (see also [22, 23, 24, 25] for non-spherical collapse of perfect fluids in asymptotically flat spacetimes).
For PBHs formed during the radiation-dominated era of the Universe, the assumption of spherical symmetry is well justified within the framework of peak theory [26]. In this context, large peaks in curvature fluctuations which are so rare that they are not overproduced beyond the dark matter density tend to be nearly spherical. However, even for such large peaks, the most likely configuration is not perfectly spherical [26]. Instead, small deviations from sphericity are typical, leading to an ellipsoidal shape characterized by parameters known as “ellipticity” ($e$) and “prolateness” ($p$) following a probability distribution. Regarding the PBH formation with the initial amplitude slightly above the threshold for which critical behavior [27] is relevant, it is not entirely clear how small deviations from sphericity impact the critical behavior of the collapse.
@@
Non-spherical effects may be relevant for some physically well-motivated scenarios where soft equations of state are considered and predictions of the mass function of PBHs may be contrasted with the gravitational wave spectrum of induced gravitational waves (see [39] for a review) and/or with merger rates of PBH binaries [40, 41]. The softening can be due to a still unknown phase transition or to the standard thermal history of the Universe [42] when particles become non-relativistic. For instance, the softening of the equation of state due to the QCD crossover about $\sim 10^{-5}{\rm s}$ after the Big Bang [43, 44, 45] might cause a significant number of PBHs in the solar mass range [46, 47, 48, 49, 50, 51, 52, 53, 54, 55]. Another possibility is from crossovers beyond standard model theories that may induce a significant softening of the equation of state [56, 57]. In addition, several scenarios with $w\neq 1/3$ have also been considered in the context of the PTA signal [58, 59] and its counterpart with PBHs [60, 61, 62, 63, 64, 65, 66].
@@ 2 Deviation from sphericity of the initial curvature perturbation in peak theory
We denote the normalized height of the peak of $\zeta$ as $\nu=\zeta(\vec{r}=0)/\sigma_{0}$ , and introduce the correlator $\gamma\equiv\sigma^{2}_{1}/(\sigma_{0}\sigma_{2})$, which depends on the power spectrum $\mathcal{P}_{\zeta}$. Let’s consider the immediate neighbour of $\zeta$ around the peak value at $\vec{r}=0$ with a Tayor expansion up through the second order as
@@
where $\lambda_{l}$ are the eigenvalues of $-\partial_{i}\partial_{j}\zeta$ and we have considered that the axes are oriented along its principal direction [^1] Even if this is not the case, so that $\zeta(\vec{r})=\zeta(\vec{r}=0)+\sum_{i}\partial_{i}\partial_{j}\zeta\frac{r_{i}r_{j}}{2}$, the symmetric coefficient matrix of the second order can always be diagonalized through a rotation transformation. with $r_{l}$ being the cartesian coordinates. We can also arrange $\lambda_{l}$, so that $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}$ by considering the rotation of the axes without loss of generality. Around the peak, the contours of $\zeta(\vec{r})$ are described by the ellipsoids given by $\zeta(\vec{r}=0)-\zeta(\vec{r})\approx\sum_{l}r^{2}_{l}/(1/\lambda_{l})={\rm const.}$ with the radius along each axis being proportional to $1/\sqrt{\lambda_{l}}$. The shape of the ellipsoid is characterized by
$e=\frac{\lambda_{1}-\lambda_{3}}{2\sum_{i}\lambda_{i}},\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,p=\frac{\lambda_{1}-2\lambda_{2}+\lambda_{3}}{2\sum_{i}\lambda_{i}}.$ (2.6)
The parameters $e$ (“ellipticity”) and $p$ (“prolateness”) are in the range $20\leq e\leq 1/$ 2 and $-e\leq p\leq e$. For instance, if $(\lambda_{1}-\lambda_{2})-(\lambda_{2}-\lambda_{3})>0\Leftrightarrow p>0$, the shape is oblate (pancake like) while it is prolate (cigar like) for $p<0$. It is also common to define “spheroids” as ellipsoids with two equal eigenvalues.
The values of $e$ and $p$ follow a specific probability distribution, which we will explore next. For instance, the probability density vanishes for $p=e$ and $p=-e$, which correspond to oblate and prolate spheroids, respectively. That is, approximately spheroidal configurations are improbable. Let’s now introduce the Gaussian distributed variable $\xi\equiv-\left.\nabla^{2}\zeta\right|_{\vec{r}=0}/\sigma_{2}=\sum_{i}\lambda_{i}/\sigma_{2}$, where we used Eq.(2.5).
@@
$+\sqrt{\frac{2}{5\pi}}\left[\mathcal{C}_{1}(\xi)\exp{-\frac{5}{8}\xi^{2}}+\mathcal{C}_{2}(\xi)\exp{-\frac{5}{2}\xi^{2}}\right]$ (2.12)
@@
$\chi&=1,\,\,\,\,0\leq e\leq 1/4,-e\leq p\leq e,\\ \chi&=1,\,\,\,\,1/4\leq e\leq 1/2,-(1-3e)\leq p\leq e,\\ \chi&=0,\,\,\,\,\textrm{otherwise.}\end{aligned}\right.$ (2.14)
This expression does not explicitly depend on $\nu$, and the domain with a non-vanishing value is restricted to |$\absolutevalue{p}\leq e$ and $e\geq 0$. Specifically, the allowed domain $(e,p)$ is the interior of a triangle bounded by the points $(0,0)$, $(1/4,-1/4)$ and $(1/2,1/2)$. The mean values of $\langle e\rangle$ and $\langle p\rangle$ with $\mathcal{P}_{e,p}$ can be computed analytically and are shown in the appendix A. For high peaks, the mean values $\langle e\rangle$ and $\langle p\rangle$ are small, and high peaks are more spherically symmetric than lower ones. Despite that, the most likely values are not exactly zero, i.e., a small deviation from sphericity likely to exist in general (see Fig.1). Notice that the mean value of $p$ is much smaller than $e$ for large $\xi$.
@@
We now focus on obtaining the typical profile of the curvature fluctuation $\bar{\zeta}$ characterized by fixed values of the parameters $(\nu,\xi,e,p)$. Euler angles are fixed, so that the coordinate axes are along the principle axes of $-\partial_{i}\partial_{j}\zeta$. Then we obtain (see (7.8) in Ref. [26])
@@
Figure 2: Top panels: Contour of $\zeta$ fixing $\zeta(x,y,z)=\mu/2$ with $\mu=0.8$ for $e=p=0$ (left panel) and $e=0.1$ and $p=0.05$ (right panel). Bottom panels: The left panel corresponds to a $3$D plot of $\zeta$ in the plane $z=0$ for $e=p=0$, and the right panel is the difference in $\zeta$ between a spherical $\zeta_{\rm sp}$ configuration and non-spherical one with $e=0.1$ and $p=0.05$.
In Appendix B, we calculate the mass function of primordial black holes (PBHs) under the assumption of spherical symmetry with a monochromatic power spectrum and different $w$’s. This calculation is later used to set a typical value for the fluctuation amplitude. We note that, the constraint from the total PBH fraction of dark matter $f_{\rm PBH}^{\rm tot}\leq 1$ requires that PBH formation is sufficiently rare, and the relevant amplitude of the peak is much larger than the standard deviation $\sigma_{0}$. That is, we need a high peak value, $\nu\gg 1$, which depends on the mass scale where $k_{p}$ is fixed (see Fig.15). The situation could differ for a broad spectrum, where $\gamma<1$, but a large peak is still likely necessary.
@@ 3 Numerical method and procedure
We perform numerical simulations following the BSSN formalism [68, 69] decomposing the spacetime metric into the following 3+1 form:
@@
where $\alpha$ is the lapse function, $\beta^{i}$ is the shift-vector, $\tilde{\gamma}_{ij}$ is the conformal metric whose determinant is given by that of the flat reference metric, $\tilde{\psi}$ is the spatial conformal factor, which at super-horizon scales is $\tilde{\psi}=a^{1/2}\Psi=a^{1/2}\exp(-\zeta/2)$ with $a$ being the scale factor of the FLRW background. All these variables are functions of $t$ and $\vec{x}$. The time evolution of these variables at super-horizon scales has been analytically investigated in Ref. [70]. The energy momentum-tensor is given by the following perfect fluid form:
@@
where $n_{\mu}=-\alpha\partial_{\mu}t$ and $\Gamma$ is the Lorentz factor defined by $\Gamma=-u^{\mu}n_{\mu}$. We use COSMOS code written in C++ [71, 72], which originally follows the SACRA code [73]. We basically use the same set-up and parameters (like the grid spacing, the CFL number condition and others) of the code used in [21] but modify the boundary conditions to handle only one quadrant in the simulations as in [18], taking into account the symmetry of our initial conditions, which makes the simulations less computationally expensive. First, we use the SPriBoSH code [74] to obtain efficiently the thresholds of PBH formation under the assumption of spherical symmetry (denoted by $\mu_{\rm c,sp}$) for several values of $w$. In this work, we consider $w=1/10$ and $1/3$ (soft and radiation cases), for what we obtain $\mu_{\rm c,sp}\approx 0.3095$ and $0.6061$, respectively (see Fig. 11 of Ref.[10] for other values of $w$). In the COSMOS code, we obtain the consistent values of $\mu_{\rm c,sp}$ for the same spherical initial profile. This also helps us to determine the optimal number of grid points needed to achieve the desired accuracy in calculating the threshold values. We have found that setting $N=60$ ($w=1/10$) and $N=100$ ($w=1/3$) grid points in each direction gives enough accuracy for our purposes. We check the correctness of our simulations using the averaged Hamiltonian constraint violation and the maximum of that quantity, for which we show detailed examples in the appendix D.
@@
Although the number of grid points is sufficient to follow the overall dynamics of the system, it is insufficient to precisely capture the apparent horizon formation without introducing the mesh-refinement procedure adopted in Ref. [21]. Instead of introducing the time-consuming mesh-refinement procedure, to determine whether a black hole forms or the fluid disperses in a FLRW background, we monitor the value of the lapse function at the origin. If the lapse function continuously decreases to a small value $\lesssim 0.1$, it suggests the eventual formation of an apparent horizon at sufficiently late times, indicating black hole formation. Conversely, if the lapse function exhibits a bouncing behaviour and increases after that, the fluid’s mass excess disperses within the FLRW background, preventing black hole formation. A similar pattern is observed in the peak value of the energy density: a continuous increase signals black hole formation while reaching a maximum followed by a decline indicates fluid dispersion. This behaviour is quite universal, appearing in both spherically symmetric and non-spherical simulations, as noted in various studies [75, 76, 74, 18, 77, 78]. In some cases, our simulation failed with a significant violation of the Hamiltonian constraint at a late time. However, the behavior of the lapse function until the significant violation of the Hamiltonian constraint allows us to reliably infer whether the fluctuation will ultimately collapse to form an apparent horizon.
@@
In the initial profile (2.16), we set $k_{p}=10L$ and $H_{0}:=H(t_{0})=50L^{-1}$ with $L$ being the coordinate length along each axis of the numerical domain given by $x\in[0,L]$, $y\in[0,L]$ and $z\in[0,L]$. The other initial quantities of the BSSN formalism are computed according to [70, 18]. The initial time corresponding to $H_{0}$ is given by $t_{0}=\alpha_{w}/H_{0}$ with $\alpha_{w}=2/(3w+3)$, and the initial energy density of the background universe is given by $\rho_{b}(t_{0})=3H^{2}_{0}/(8\pi)$. We also define the time of horizon crossing $t_{H}$ of the mode $k_{p}$ as $t_{H}=t_{0}(H_{0}/k_{p})^{1/(1-\alpha_{w})}$.
@@
$W(r)=\exp{-\left(\frac{r-r_{W}}{\sigma_{W}}\right)^{\alpha_{W}}}$ ) (3.4)
with values $\sigma_{W}=0.09L$, $r_{W}\approx 0.77L$ and $\alpha_{W}=4$. Effectively, the window function will change the functional form of Eq.(2.16), but only for a large radius $r>r_{W}$. As found in [28], the threshold of formation is mainly affected by the shape around the radius at which the compaction function takes the maximum value. [^3] The definition of the compaction function, which is defined for spherically symmetric fluctuations goes beyond the scope of the paper. We refer the reader to [79] for the original definition. This radius is much smaller than the radius where the window function is introduced. This expectation is confirmed in spherical symmetry by comparing the threshold with that obtained by the spherically symmetric code SPriBoSH without introducing the window function. The difference in the threshold values is within $0.01\%$. Although our case is now a simulation beyond spherical symmetry, we expect such consideration to also hold [^4] The situation would differ for the behaviour of the PBH mass, where the shape of the profile beyond the maximum of the compaction function may have a significant effect [80, 81]..
@@
Table 1: Parametters $\mu_{c}$, $\mu_{t}$, $\sigma_{0}$, $\nu_{c}$ satisfying $f_{\rm PBH}^{\rm tot}\simeq 1$ for both cases of $w$ with $k_{p}=10^{13.5}{\rm Mpc^{-1}}$ (asteroid mass range) and $k_{p}=10^{7}{\rm Mpc^{-1}}$ (solar mass range).

### v3

@@ 1 Introduction
Primordial black holes (PBHs), if they exist, are black holes that could have been formed in the early Universe [1, 2, 3, 4, 5] (see [6, 7, 8, 9, 10, 11] for reviews covering different perspectives). A PBH results from the collapse of a sufficiently large over-density in the early Universe, much before the moment of matter-radiation equality, and therefore, their existence can be evidence of the existence of primordial inhomogeneities. Interestingly, PBHs may constitute a significant fraction of the dark matter [12] (especially in the so-called asteroid mass range) and explain different cosmic conundra [13]. Currently, PBHs have not been detected so far, but future gravitational wave observations may establish their existence (in particular if their mass is lower than a solar mass)[14, 15, 16] and quantify their role in the dark matter.
@@
To accurately estimate the abundance of PBHs in our Universe, it is essential to clarify the initial conditions that lead to their formation. Particularly, determinating the threshold is essential for the abundance estimation as it is exponentially sensitive to the threshold [4]. The threshold estimation typically requires relativistic numerical simulations. Most studies on simulation of PBH formation have focused on spherically symmetric systems (see [17] for a comprehensive review and references therein) for which we have nowadays a relatively solid understanding. In contrast, our understanding of PBH formation in non-spherical scenarios currently remains limited, with only a few studies numerically addressing this issue [18, 19, 20, 21] (see also [22, 23, 24, 25] for non-spherical collapse of perfect fluids in asymptotically flat spacetimes).
For PBHs formed during the radiation-dominated era of the Universe, the assumption of spherical symmetry is well justified within the framework of peak theory [26]. In this context, large peaks in curvature fluctuations which are so rare that they are not overproduced beyond the dark matter density tend to be nearly spherical. However, even for such large peaks, the most likely configuration is not perfectly spherical [26]. Instead, small deviations from sphericity are typical, leading to an ellipsoidal shape characterized by parameters known as “ellipticity” ($e$) and “prolateness” ($p$) following a probability distribution. Regarding the PBH formation with the initial amplitude slightly above the threshold for which critical behavior [27] is relevant, it is not entirely clear how small deviations from sphericity impact the critical behavior of the collapse. Therefore, it is crucial to carefully examine how even a small deviation from sphericity might influence gravitational collapse and the corresponding statistical estimates of the PBH mass function and abundance. This analysis is necessary to contrast with the typical assumptions in the literature, where spherical symmetry is often assumed. For this analysis, relativistic $3+1$ simulations are essential.
@@
Non-spherical effects may be relevant for some physically well-motivated scenarios where soft equations of state are considered and predictions of the mass function of PBHs may be contrasted with the gravitational wave spectrum of induced gravitational waves (see [39] for a review) and/or with merger rates of PBH binaries [40, 41]. The softening can be due to a still unknown phase transition or to the standard thermal history of the Universe [42] when particles become non-relativistic. For instance, the softening of the equation of state due to the QCD crossover about $\sim 10^{-5}{\rm s}$ after the Big Bang [43, 44, 45] might cause a significant number of PBHs in the solar mass range [46, 47, 48, 49, 50, 51, 52, 53, 54, 55]. Another possibility is from crossovers beyond standard model theories that may induce a significant softening of the equation of state [56, 57]. In addition, several scenarios with $w\neq 1/3$ have also been considered in the context of the pulsar timing array (PTA) signal [58, 59] and its counterpart with PBHs [60, 61, 62, 63, 64, 65, 66].
@@ 2 Deviation from sphericity of the initial curvature perturbation in peak theory
We denote the normalized height of the peak of $\zeta$ as $\nu=\zeta(\vec{r}=0)/\sigma_{0}$ , and introduce the correlator $\gamma\equiv\sigma^{2}_{1}/(\sigma_{0}\sigma_{2})$, which depends on the power spectrum $\mathcal{P}_{\zeta}$. Let’s consider the immediate neighbour of $\zeta$ around the peak value at $\vec{r}=0$ with a Taylor expansion up through the second order as
@@
where $\lambda_{l}$ are the eigenvalues of $\partial_{i}\partial_{j}\zeta$ and we have considered that the axes are oriented along its principal direction [^1] Even if this is not the case, so that $\zeta(\vec{r})=\zeta(\vec{r}=0)+\sum_{i}\partial_{i}\partial_{j}\zeta\frac{r_{i}r_{j}}{2}$, the symmetric coefficient matrix of the second order can always be diagonalized through a rotation transformation. with $r_{l}$ being the Cartesian coordinates. We can also arrange $\lambda_{l}$, so that $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}$ by considering the rotation of the axes without loss of generality. Around the peak, the contours of $\zeta(\vec{r})$ are described by the ellipsoids given by $\zeta(\vec{r}=0)-\zeta(\vec{r})\approx\sum_{l}r^{2}_{l}/(1/\lambda_{l})={\rm const.}$ with the radius along each axis being proportional to $1/\sqrt{\lambda_{l}}$. The shape of the ellipsoid is characterized by
$e=\frac{\lambda_{1}-\lambda_{3}}{2\sum_{i}\lambda_{i}},\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,p=\frac{\lambda_{1}-2\lambda_{2}+\lambda_{3}}{2\sum_{i}\lambda_{i}}.\$(2.6)
The parameters $e$ (“ellipticity”) and $p$ (“prolateness”) are in the range $0\leq e\leq 1/2$ and $e\leq p\leq e$. For instance, if $(\lambda_{1}-\lambda_{2})-(\lambda_{2}-\lambda_{3})>0\Leftrightarrow p>0$, the shape is oblate (pancake-like) while it is prolate (cigar-like) for $p<0$. It is also common to define “spheroids” as ellipsoids with two equal eigenvalues.
The values of $e$ and $p$ follow a specific probability distribution, which we will explore next. For instance, the probability density vanishes for $p=e$ and p=ep=-e, which correspond to oblate and prolate spheroids, respectively. That is, approximately spheroidal configurations are improbable. Let’s now introduce the Gaussian distributed variable $\xi\equiv-\left.\nabla^{2}\zeta\right|_{\vec{r}=0}/\sigma_{2}=\sum_{i}\lambda_{i}/\sigma_{2}$, where we used Eq.(2.5).
@@
[…]
@@
[…]
@@
We now focus on obtaining the typical profile of the curvature fluctuation $\bar{\zeta}$ characterized by fixed values of the parameters $(\nu,\xi,e,p)$. Euler angles are fixed, so that the coordinate axes are along the principle axes of $\partial_{i}\partial_{j}\zeta$. Then we obtain (see (7.8) in Ref. [26])
@@
Figure 2: Top panels: Contour of $\zeta$ fixing $\zeta(x,y,z)=\mu/2$ with $\mu=0.8$ for $e=p=0$ (left panel) and $e=0.1$ and $p=0.05$ (right panel). Bottom panels: The left panel corresponds to a 33D plot of $\zeta$ in the plane $z=0$ for $e=p=0$, and the right panel is the difference in $\zeta$ between a spherical $\zeta_{\rm sp}$ configuration and non-spherical one with $e=0.1$ and $p=0.05$.
In Appendix B, we calculate the mass function of primordial black holes under the assumption of spherical symmetry with a monochromatic power spectrum and different $w$’s. This calculation is later used to set a typical value for the fluctuation amplitude. We note that, the constraint from the total PBH fraction of dark matter $f_{\rm PBH}^{\rm tot}\leq 1$ requires that PBH formation is sufficiently rare, and the relevant amplitude of the peak is much larger than the standard deviation $\sigma_{0}$. That is, we need a high peak value, $\nu\gg 1$, which depends on the mass scale where $k_{p}$ is fixed (see Fig.15). The situation could differ for a broad spectrum, where $\gamma<1$, but a large peak is still likely necessary.
@@ 3 Numerical method and procedure
We perform numerical simulations following the Baumgarte–Shapiro–Shibata–Nakamura (BSSN) formalism [68, 69] decomposing the spacetime metric into the following 3+1 form:
@@
where $\alpha$ is the lapse function, $\beta^{i}$ is the shift vector, $\tilde{\gamma}_{ij}$ is the conformal metric whose determinant is given by that of the flat reference metric, and $\tilde{\psi}$ is the spatial conformal factor, which at super-horizon scales is $\tilde{\psi}=a^{1/2}\Psi=a^{1/2}\exp(-\zeta/2)$ with $a$ being the scale factor of the Friedmann-Lemaître-Robertson-Walker (FLRW) background. All these variables are functions of $t$ and $\vec{x}$. The time evolution of these variables at super-horizon scales has been analytically investigated in Ref. [70]. The energy momentum-tensor is given by the following perfect fluid form:
@@
where $n_{\mu}=-\alpha\partial_{\mu}t$ and $\Gamma$ is the Lorentz factor defined by $\Gamma=-u^{\mu}n_{\mu}$. We use COSMOS code written in C++ [71, 72], which originally follows the SACRA code [73]. The code uses the mono upstream-centered scheme for conservation laws (MUSCL) [74, 75] for the fluid dynamics (we refer the reader to Appendix E for a summary of the evolution scheme used). We basically use the same setup and parameters (like the grid spacing, the Courant–Friedrichs–Lewy number condition and others) of the code used in [21] but modify the boundary conditions to handle only one quadrant in the simulations as in [18], taking into account the symmetry of our initial conditions, which makes the simulations less computationally expensive. First, we use the SPriBoSH code [76] to obtain efficiently the thresholds of PBH formation under the assumption of spherical symmetry (denoted by $\mu_{\rm c,sp}$) for several values of $w$. In this work, we consider $w=1/10$ and $1/3$ (soft and radiation cases), for what we obtain $\mu_{\rm c,sp}\approx 0.3095$ and $0.6061$, respectively (see Fig. 11 of Ref.[10] for other values of $w$). In the COSMOS code, we obtain the consistent values of $\mu_{\rm c,sp}$ for the same spherical initial profile. This also helps us to determine the optimal number of grid points needed to achieve the desired accuracy in calculating the threshold values and make a control test. We have found that setting $N=60$ ($w=1/10$) and $N=100$ ($w=1/3$) grid points in each direction gives enough accuracy for our purposes. We check the correctness of our simulations using the averaged Hamiltonian constraint violation, for which we show detailed examples in the Appendix D.
@@
Although the number of grid points is sufficient to follow the overall dynamics of the system, it is insufficient in some cases to precisely capture the apparent horizon formation without introducing the mesh-refinement procedure adopted in Ref. [21]. Instead of introducing the time-consuming mesh-refinement procedure, to determine whether a black hole forms or the fluid disperses in a FLRW background, we monitor the value of the lapse function at the origin. If the lapse function continuously decreases to a small value $\lesssim 0.1$, it suggests the eventual formation of an apparent horizon at sufficiently late times, indicating black hole formation. Conversely, if the lapse function exhibits a bouncing behaviour and increases after that, the fluid’s mass excess disperses within the FLRW background, preventing black hole formation. A similar pattern is observed in the peak value of the energy density: a continuous increase signals black hole formation while reaching a maximum followed by a decline indicates fluid dispersion, although observing the full process of dispersion of the fluctuation requires much more computational time than what we have used. This behaviour is quite universal, appearing in both spherically symmetric and non-spherical simulations, as noted in various studies [77, 78, 76, 18, 79, 80]. In some cases, our simulation failed with a significant violation of the Hamiltonian constraint at a late time. However, the behavior of the lapse function until the significant violation of the Hamiltonian constraint allows us to reliably infer whether the fluctuation will ultimately collapse to form an apparent horizon.
@@
In the initial profile (2.16), we set $k_{p}=10L^{-1}$ and $H_{0}:=H(t_{0})=50L^{-1}$ with $L$ being the coordinate length along each axis of the numerical domain given by $x\in[0,L]$, $y\in[0,L]$ and $z\in[0,L]$. The other initial quantities of the BSSN formalism are computed according to [70, 18]. The initial time corresponding to $H_{0}$ is given by $t_{0}=\alpha_{w}/H_{0}$ with $\alpha_{w}=2/(3w+3)$, and the initial energy density of the background Universe is given by $\rho_{b}(t_{0})=3H^{2}_{0}/(8\pi)$. We also define the time of horizon crossing $t_{H}$ of the mode $k_{p}$ as $t_{H}=t_{0}(H_{0}/k_{p})^{1/(1-\alpha_{w})}$.
@@
[…]
with values $\sigma_{W}=0.09L$, $r_{W}\approx 0.77L$ and $\alpha_{W}=4$. Effectively, the window function will change the functional form of Eq.(2.16), but only for a large radius $r>r_{W}$. As found in [28], the threshold of formation is mainly affected by the shape around the radius at which the compaction function takes the maximum value. [^3] The definition of the compaction function, which is defined for spherically symmetric fluctuations goes beyond the scope of the paper. We refer the reader to [81] for the original definition. This radius is much smaller than the radius where the window function is introduced. This expectation is confirmed in spherical symmetry by comparing the threshold with that obtained by the spherically symmetric code SPriBHoS without introducing the window function. The difference in the threshold values is within $0.01\%$. Although our case is now a simulation beyond spherical symmetry, we expect such consideration to also hold [^4] The situation would differ for the behaviour of the PBH mass, where the shape of the profile beyond the maximum of the compaction function may have a significant effect [82, 83]..
@@
Table 1: Parameters $\mu_{c}$, $\mu_{t}$, $\sigma_{0}$, $\nu_{c}$ satisfying $f_{\rm PBH}^{\rm tot}\simeq 1$ for both cases of $w$ with $k_{p}=10^{13.5}{\rm Mpc^{-1}}$ (asteroid mass range) and $k_{p}=10^{7}{\rm Mpc^{-1}}$ (solar mass range).